et us say the income elasticity of money demand is 0.5 and the interest elasticity of money demand is –0.3 (as the empirical studies suggest). What is the percentage change of the real money demand if income increases 5.5% at the same time that the interest rate decreases from 4.2% to 3.8%?
1
Expert's answer
2020-06-12T12:01:28-0400

"\frac{M}{P} = k\times Y-h\times i"


"\frac{M}{P} =105.5\times0.5 Y-(-0.3\times 3.8)-=0.5275+1.14=1.6675"